OpenStudy (anonymous):

An unfair coin (0.4 probability for heads, 0.6 probability for tails) is tossed 5 times. a) What is the probability that you get exactly 3 tails? b) What is the probability that you get at least 3 tails?

OpenStudy (vivek3461):

Probability of getting exactly 3 tails = 0.6*0.6*0.6*0.4*0.4 No of possible combinations for getting exactly 3 tails = 5C3 So the total probability is 5C3*0.6*0.6*0.6*0.4*0.4

OpenStudy (anonymous):

I am sorry. 5C3? :) Thank you very much. I am new with statistics.

OpenStudy (anonymous):

\[_5 C_3~means~5~Combination~3...~written~in~formula~as...\]\[_5 C_3=\frac{n!}{r!(n-r)!}=\frac{5!}{3!2!}=\frac{(5)(4)(3)(2)(1)}{(3)(2)(1)(2)(1)}=\frac{(5)(4)}{2}=10\]